What is the pH of 40.0 mL of a solution that is 0.17 M in CN and 0.25 M in HCN For HCN, use K-49-10-9. Number
Answer #1

PKa = -log(Ka)

= -log(4.9*10^-9)

=8.31

pH = PKa + log (CN- / HCN)

pH = 8.31 + log (0.17 / 0.25)

8.143

any query please comment

if satisfied please rate